# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v1

@@ II-D2 Information Retrieval
Given queries and documents, the IR model can provide their matching scores, and the quality of the model can be evaluated through these scores. For the first stage, common evaluation metrics include $\text{Recall}@k$ and $\text{Accuracy}@k$, while for the second stage, metrics like MAP, MRR and NDCG are generally used.
$\text{Recall}@k$ calculates a truncated Recall value at the $k$-th position of a retrieved list.
$\text{Recall}@k=\frac{1}{\left|\mathcal{Q}\right|}\sum_{q=1}^{\left|\mathcal{Q}\right|}\frac{R_{q,k}}{S_{q}},$ (4)
where $\mathcal{Q}$ is the query set. $S_{q}$ denotes the total number of relevant texts for query q, and $R_{q,k}$ denotes the number of relevant texts retrieved in the top-$k$ positions by the retrieval model.
$\text{Accuracy}@k$ calculates the proportion of queries for which the top-$k$ retrieved texts contain the answers, defined as:
$\text{Accuracy}@k=\frac{1}{\left|\mathcal{Q}\right|}\sum_{q=1}^{\left|\mathcal{Q}\right|}\mathbb{I}(R_{q,k}>0),$ (5)
where $\mathbb{I}(\cdot)$ is an binary indicator function that only returns 1 when the case is true.
Furthermore, Mean Average Precision (MAP) calculates the mean of Precision scores over a set of queries $\mathcal{Q}$:
$\text{MAP}=\frac{1}{\left|\mathcal{Q}\right|}\sum_{q=1}^{\left|\mathcal{Q}\right|}\text{AP}_{q},$ (6)
where $\text{AP}_{q}=\frac{1}{S_{q}}\sum_{k=1}^{L}\text{Precision}@k\times\mathbb{I}(q,k)$, and $\text{Precision}@k=\frac{1}{\left|\mathcal{Q}\right|}\sum_{q=1}^{\left|\mathcal{Q}\right|}\frac{R_{q,k}}{k}$.
MRR, averages the reciprocal of the rank of the first retrieved positive text over a set of queries $\mathcal{Q}$
$\text{MRR}=\frac{1}{\left|\mathcal{Q}\right|}\sum_{q=1}^{\left|\mathcal{Q}\right|}\frac{1}{\text{rank}_{q}},$ (7)
where $\text{rank}_{q}$ is the position of the first retrieved positive text.
In addition to quantifying the frequency of positive texts, Normalized Discounted Cumulative Gain (DCG) also considers the position of a relevant text. It favors a ranking system that situates a pertinent text at a superior position.
$\text{DCG}_{q}@k=\sum_{i=1}^{k}\frac{2^{r_{i}}-1}{\log_{2}(i+1)},$ (8)
where $r_{i}$ is the graded relevance score for the $i$-th retrieved text. According to the aforementioned definition of DCG, NDCG is the aggregate of the normalized DCG values at a specific rank position.
$\text{NDCG}@k=\frac{1}{\left|\mathcal{Q}\right|}\sum_{q=1}^{\left|\mathcal{Q}\right|}\frac{\text{DCG}_{q}@k}{\text{IDCG}_{q}@k},$ (9)
where IDCG@$k$ denote ideal discounted cumulative gain at a particular rank position $k$.
II-D3 Comprehensive Benchmark
Although different downstream tasks require different features to be included in text embeddings, training a model to get a more general embedding for all texts has been a hot research topic. The MTEB benchmark [72] contains 56 datasets in 7 different types of downstream tasks, including classification, clustering, pair classification, reranking, retrieval, and Semantic Textual Similarity (STS), to assess the degree of generalization of text embedding. Until now, MTEB has been available in Chinese [73], French [74], Polish [75], Scandinavian [76], Russian [77] and Arabic [78].
II-E Emerging Related Tasks
II-E1 Long Context Compression
Long Context Compression involves compressing long context tokens for large models without sacrificing essential information, accelerating inference while maintaining performance in scenarios with extensive token lengths.
@@ II-E1 Long Context Compression
The Question Answer (QA) dataset for Long Context Compression task can be formalized as follows: $D=\{(C_{i},q_{i},a_{i})\}_{i=1}^{n}$ where $C_{i}$ represents the long context, $q_{i}$ represents the question, $a_{i}$ represents the correct answer to the question $q_{i}$, derived from the context $C_{i}$. The training aims to allow the model to extract the correct answers from both the original long context $C_{i}$ and the compressed context $C^{\prime}_{i}$. Evaluation is performed by measuring the model’s ability to provide the correct answer $a_{i}$ even when the context has been compressed.
@@
In Long Context Compression, evaluations are generally divided into two approaches. The first is to assess compression efficiency using metrics like context compression rate [79] and average inference time [80]. The second is to directly evaluate the generated quality by applying downstream QA tasks. In QA tasks, the generated answers are typically compared to the reference answers using various metrics to calculate consistency scores. Common metrics include Perplexity [81], Exact Match [80], and ROUGE-L [80].
@@
Context Compression Rate indicates The ratio of the compressed context length to the original context length:
$\text{CompressionRate}=\frac{\text{len}(C^{\prime})}{\text{len}(C)}$ (10)
@@
Perplexity measures how well the model predicts the next word in a sequence, indicating its confidence and fluency:
$\text{Perplexity}=(\prod_{i=1}^{n}\frac{1}{p(w_{i}|w_{1},...,w_{i-1})})^{\frac{1}{n}}$ (11)
@@
$\text{EM}=\frac{\sum_{i=1}^{N}\mathbb{I}(\text{prediction}_{i}=\text{groundtruth}_{i})}{N}$ (12)
@@
II-E2 Embedding Inversion
Embedding inversion is a type of privacy leakage that predicts relevant information based on the embedding of the input text. It is generally regarded as an attacker, with targets including attribute information in the input text, the entire input text, or the text embedding model itself.
@@ II-E2 Embedding Inversion
For different levels of the leak information, the datasets can be word-level or sentence-level. For attribute inference attack, the dataset has the form of (Input,Words), where the ”Input” is text and ”words” represents important information in the input text, for example name, ID number. For the purpose of reconstruct all the input text, the dataset just contains input text. It should be noted that, text embedding will not be contained in dataset directly, it is generated by the victim embedding model using the text in dataset as input.
@@
The main metrics used for evaluating embedding inversion include BLEU, ROUGE, Exact-Match, cosine similarity, and Token F1.
BLEU is used to measure n-gram similarity between reconstructed text and original input text. we can compute it following:
@@
where BP is the brevity penality, $w_{n}$ is the weight for each n-gram precision and $p_{n}$ is the modified precision for n-th n-gram.
ROUGE represents the recall of overlapping words of reconstructed text and Exact-Match represents the percentage of exactly matching reconstructed text to the original input text. While the cosine similarity is computed in the embedding space between embeddings of reconstructed text and original input text. The ”Token F1” metric is the multi-class F1 scores between tokens of predicted text and input text and It is word-level metric.
TABLE III: The Overview of LLM-augmented Text Embedding. I, Q, D+, D- denote instruction, query, positive document, and negative document separately.
@@
AnglE [85] ChatGPT, LLaMA2, ChatGLM BERT 6K (Q, D+, D-) $\surd$ ZS STS
@@
SumCSE [87] Vicuna RoBERTa 276K (Q, D+, D-) $\times$ ZS STS
@@
SKICSE [89] LLaMA2 BERT 1M, 276K (Q, D+) $\times$ ZS STS
DenoSent [90] ChatGPT BERT, RoBERTa 1M (Q, D+) $\surd$ ZS STS
@@
InPars [94] ChatGPT monoT5 10K (Q, D+) $\surd$ ZS / FS IR
InPars-v2 [95] GPT-J monoT5 180K (Q, D+) $\surd$ ZS / FS IR
I3 [96] ChatGPT COCO-DR(BERT) 140K (I, Q, D+) $\times$ ZS IR
@@
NV-Retriever [99] E5mistral Mistral 956K (Q, D-) $\times$ ZS / FT IR
@@
E5mistral [102] ChatGPT, GPT4 Mistral 500K (I, Q, D+) $\times$ ZS Uni.
@@
III LLM-Augmented Text Embedding
One approach to adopting LLMs for text embeddings is through knowledge distillation. Specifically, LLMs can be used to generate training data for the embedding model in two ways: 1. Directly synthesizing training data (§ III-A); 2. Providing supervision signals for existing data (§ III-B). A detailed description of different methods is presented in Table III.
III-A Data Synthesis with LLMs
@@
III-A1 Instructions
Existing instructions in training data are typically constructed based on datasets. This involves manually collecting multi-task datasets and applying templates to generate different instructions for each dataset [104]. However, this approach lacks diversity, limiting the full potential of the model’s instruction-following capabilities. Therefore, a promising solution is to generate instance-level instructions by using LLMs. One approach directly leverages LLMs to generate instructions: I3 [96] generates diverse instructions by manually setting different conditions, including topics, organizational formats of the retrieved text, and definitions of relevance; E5mistral [102] categorizes tasks into different types (e.g., symmetric and asymmetric retrieval tasks) and designs specific prompts for each category to generate instructions. Another approach generates instructions based on existing data: Gecko[101] prompts LLMs to generate instructions by conditioning on different documents; Promptriever[100] generates instructions that describe the relationship between the given queries and documents, and enhance instruction diversity by setting various conditions such as length and style of the generated instructions.
III-A2 Positive Samples
@@
A classic task where anchors and positive samples are symmetric is STS. In this area, numerous studies have proposed various methods to generate symmetric positive samples. Based on the analysis in S-BERT [105], text embedding models for STS are often trained using training data from the Natural Language Inference (NLI) task. Following this convention, a line of research explores how to use LLMs to generate NLI-style training data: Two studies [86, 92] directly utilize the in-context learning capabilities of LLMs by using NLI-formatted training data as demonstrations to guide the generation of entailment sentences as positive samples; Other studies [82, 84, 88] first fine-tune LLMs with NLI training data, enabling them to generate NLI-formatted positive samples. Notably, AdaptCL [88] employ a reinforcement learning approach to iteratively update both the data generation model and the text embedding model, achieving improved performance. In addition, some studies explore generating other types of semantically similar texts as positive samples: Two studies [83, 85] generate diverse semantically similar positive samples by setting different conditions (e.g., genre, topic) and leveraging the in-context learning capabilities of LLMs; SKICSE [89] constructs positive samples by extracting additional information about the anchor using the inherent knowledge of LLMs. SumCSE [87] generates positive samples by having LLMs summarize the anchor; CLAIF [91] creates positive samples by masking tokens in the anchor and completing the missing parts using LLMs; E5mistral [102] further subdivides symmetric retrieval tasks into STS tasks and bitext retrieval tasks, designing specific prompts to generate different types of positive samples for each.
@@
In asymmetric cases, the anchor and positive sample are often referred to as the query and document, respectively. In various studies, either the query or the document can be the target for generation by LLM s. Many works [93, 94, 95, 97, 98, 101] use the vast amount of documents available on the internet to generate queries: Some of them [93, 94, 95, 97, 101] directly leverage the in-context learning capability of LLMs, constructing prompts to facilitate query generation. Notably, InPars [94, 95] introduces a “Guided by Bad Questions (GBQ)” template, which uses randomly sampled pseudo-queries to improve the quality of generated queries; Moreover, another study [98] employs soft prompts to fine-tune LLMs, enabling them to generate queries based on documents. In addition to generating queries from documents, there are other generation methods: I3 [96] and E5mistral [102] generate both queries and documents based on their self-generated instructions; FollowIR [103] and promptriever [100] guides document generation using queries and the generated instructions.
III-A3 Negative Samples
Negative samples are crucial in contrastive learning, and research [106] has shown that hard negatives, as opposed to randomly sampled negatives from training batches, can significantly improve model performance. Consequently, many studies have explored methods for generating hard negative samples using LLMs. When generating hard negatives, methods similar to those for generating positive samples (as discussed in § III-A2) are often employed [82, 84, 86, 88, 92, 83, 85, 87, 100]. The generated hard negatives can generally be divided into two categories. 1. NLI-based negatives [82, 84, 86, 88, 92]: Texts with a logical relationship of contradiction are generated as hard negatives; 2. Contextually irrelevant negatives [83, 85, 87, 100]: Texts are generated using specific definitions of irrelevance as hard negatives. Notably, Promptriever [100] introduces a unique type of hard negative to enhance the model’s instruction-following capability. These hard negatives are positive samples relative to the query itself but become negative samples when the query is combined with the instruction.
III-B Data Annotation with LLMs

### v2

@@ 2.3.1 Semantic Textual Similarity
For evaluation, all texts $\{x_{i}^{a}\}_{i=1}^{n}\cup\{x_{i}^{b}\}_{i=1}^{n}$ are encoded by the learned embedder. And a similarity function, e.g. dot product or cosine similarity, is introduce to measure the semantic similarity between the text pairs in the embedding space. Common evaluation metrics include Pearson correlation coefficient and Spearman correlation coefficient.
  • Pearson correlation coefficient primarily assesses the correlation between the ground-truth $c_{i}$ and the predicted similarity $c_{i}^{p}$, which is expressed as
    where $\bar{c}_{i}$ and $\bar{c}_{i}^{p}$ are the mean value of $\{c_{i}\}_{i=1}^{n}$ and $\{c_{i}^{p}\}_{i=1}^{n}$, respectively.
  • Spearman correlation coefficient primarily assesses the correlation between the ranks of $c_{i}$ and $c_{i}^{p}$ in their respective lists, which is expressed as
    $r_{\rm spearman}=1-\frac{6\sum_{i=1}^{n}(r_{i}-r_{i}^{p})^{2}}{n(n^{2}-1)}$ (3)
    where $r_{i}$ and $r_{i}^{p}$ are the rank of $c_{i}$ in $\{c_{i}\}_{i=1}^{n}$ and $c_{i}^{p}$ in $\{c_{i}^{p}\}_{i=1}^{n}$, respectively.
Since different methods lead to large differences in the value range of similarity, and we are concerned with relative rankings rather than absolute values in practice [118]. Therefore, Spearman correlation coefficient is used by default.
2.3.2 Information Retrieval
Information retrieval (IR)  [^7] Information retrieval in this survey refers to ad-hoc text retrieval by default. aims to retrieve the most related texts relevant to the query from a large-size candidate sets. Modern IR systems involves multiple stages of recalling or re-ranking from different-scale candidate sets. Embedding-based retrieval methods are referred to as dense retrieval, which is often used as a method of recalling or secondary-filtering, placed after sparse retrieval (e.g., BM25 [119]) and before reranking (e.g., the cross-encoder scheme recommended by BERT  [68]).
@@ 2.3.2 Information Retrieval
The dataset for evaluating dense retrieval can be written as $D_{\rm eval}^{\rm IR}=\{q_{i},d_{i}^{+}\}_{i=1}^{n}\cup\{d_{j}^{-}\}_{j=1}^{N}$, where $q_{i}$ is the query, $d_{i}^{+}$ are the related document of $q_{i}$ and $d_{j}^{-}$ is the unrelated text of any $q_{i}$. In the dataset, $N$ is usually much larger than $n$. The popular benchmarks include BEIR [120] for the English setting, MIRACL [57] and Mr.TyDi [58] for the multi-lingual setting.
@@
For efficiency of evaluation, all query $\{q_{i}\}_{i=1}^{n}$ and all candidate texts $\{d_{i}\}_{i=1}^{n}\cup\{d_{j}\}_{j=1}^{N}$ can be encoded as embeddings with the learned embedder in advance. Then a similarity function is introduced to measure the relevance for each query-candidate embedding pair. For each query, all candidate texts will sorted by the predicted relevance in descending and form a ordered candidate list. The common evaluation metrics include Recall Rate, Accuracy, Mean Reciprocal Rank (MRR), Mean Average Precision (MAP) and Normalized Discounted Cumulative Gain (NDCG).
  • Recall Rate calculates a truncated recall value at the $k$-th position of a sorted candidate list.
    $\text{Recall}@k=\frac{T_{q_{i}}^{(k)}}{S_{q_{i}}}$ (4)
    where $S_{q_{i}}$ denotes the total number of relevant texts for query $q_{i}$, and $T_{q_{i}}^{(k)}$ denotes the relevant text number in the top-$k$ position of the sorted candidate list.
  • Accuracy calculates the proportion of queries for which the top-$k$ retrieved texts contain the answers, defined as
    $\text{Accuracy}@k=\mathbb{I}(T_{q_{i}}^{(k)}>0),$ (5)
    where $\mathbb{I}(\cdot)$ is an indicator function.
  • MRR calculates the average reciprocal rank of the first retrieved related text over all queries, which can be denoted as
    $\text{MRR}=\frac{1}{\text{R}_{q_{i}}^{(1)}}$ (6)
    where $\text{R}_{q_{i}}^{(1)}$ is the rank of the frist relevant text in the ordered candidate list.
  • MAP calculates the mean value of Precision over all queries, which can be expressed as
    $\text{MAP}=\frac{1}{S_{q_{i}}}\sum_{k=1}^{|S_{q_{i}}|}\text{Precision}@R_{q_{i}}^{(k)}$ (7)
    where $\text{R}_{q_{i}}^{(k)}$ is the rank of the $k$-th relevant text in the sorted candidate list and $\text{Precision}@k=T_{q_{i}}^{(k)}/k$.
  • NDCG considers the rank of the relevant text, and suggest situating the more pertinent text at a more superior position. DCG for each $q_{i}$ should be calculated first, which can be written as
    $\text{DCG}_{q_{i}}@k=\sum_{i=1}^{k}\frac{2^{r_{i}}-1}{\log_{2}(i+1)}$ (8)
    where $r_{i}$ is the graded relevance score for the $i$-th retrieved text in the candidae text list. Then $\text{NDCG}@k$ can be calculated by aggregating the normalized DCG values at a specific rank position:
    $\text{NDCG}@k=\frac{\text{DCG}_{q}@k}{\text{IDCG}_{q}@k}$ (9)
    where IDCG@$k$ denote ideal DCG (the DCG value when assuming that the retrieved results are ordered optimally).
2.3.3 Universal Embedding
While tasks such as classification, rerank, summarization, etc., can be solved using the LLMs’ generative paradigm, the state-of-the-art methods require complex handwritten prompts, many in-context examples and long output for chain-of-thought. Using LLM embedding to solve these tasks is still an efficient and practical approach [121]. Therefore, training an embedder to get a more general embedding for all tasks is also a hot research topic.
Evaluation Dataset
The MTEB benchmark [122] contains 56 datasets in 7 different types of downstream tasks, including classification, clustering, pair classification, reranking, retrieval, and Semantic Textual Similarity (STS), to assess the degree of generalization of text embedding. Until October 2024, MTEB has the community variants in Chinese [123], French [124], Polish [125], Scandinavian [126], Russian [127], Arabic [128] and Persian [129], and a official multilingual version, i.e. MMTEB  [^8] https://github.com/embeddings-benchmark/mteb/blob/main/docs/mmteb, during developing.
Evaluation Metric
MTEB has developed a uniform evaluation protocol for each type of task for a fair comparison. Please refer to their original paper [122] and official github repository  [^9] https://github.com/embeddings-benchmark/mteb for the details of the evaluation metric.
2.4 Emerging Related Tasks
2.4.1 Long Context Compression
Long Context Compression (ICC) aims to compress long context into text embedding or token sequence for LLMs without sacrificing essential information, accelerating inference while maintaining generation results. LLMs are expected to be interpreters that understand original context information from compressed embedding or token sequence and output content consistent with what it would output when there is no compression.
Evaluation Dataset
The evaluation for LCC usually uses the dataset for Question Answering (QA) or Retrieval Augmented Generation (RAG). The datasets are formalized as $D_{\rm LCC}^{\rm eval}=\{(c_{i},q_{i},a_{i})\}_{i=1}^{n}$ where $c_{i}$ represents the $L_{c_{i}}$-token context, $q_{i}$ represents the question, and $a_{i}=\{t_{i}^{(j)}\}_{j=1}^{L_{a_{i}}}$ represents the $L_{a_{i}}$-token answer.
Evaluation Metric
The embedder for LCC takes the L-length long context $c_{i}$ as input and outputs a compressed embedding matrix $c_{i}^{\prime}\in\mathbb{R}^{d\times l_{i}}$. The evaluation for LCC can be divided into two aspects: (1) compression efficiency and (2) generation consistency. Specifically, context compression rate [130] and average inference time [131] are proposed to evaluate the compression efficiency, while Perplexity [132], Exact Match [131], and ROUGE-L [131] are suggested for the evaluation of generation consistency.
@@ 2.4.1 Long Context Compression
  • Context Compression Rate measures the ratio of the compressed context length to the original context length, which is denoted as
    $\text{CompressionRate}=\frac{l_{i}}{L_{c_{i}}}$ (10)
@@
  • Perplexity measures how well the model predicts the next word in the answer, indicating its confidence and fluency:
    $\text{Perplexity}=\prod_{i=1}^{L_{a_{i}}}\frac{1}{p(t_{i}^{(j)}|t_{i}^{(1)},...,t_{i}^{(j-1)})})^{\frac{1}{L_{a_{i}}}}$ (11)
@@
    $\text{EM}=\frac{\sum_{i=1}^{\operatorname{min}\{L_{a_{i}},L_{\hat{a}_{i}}\}}\mathbb{I}(\hat{t}_{a_{i}}^{(j)}=t_{i}^{(j)})}{L_{a_{i}}}$ (12)
    where $\hat{t}_{a_{i}}^{(j)}$ is the $j$-th token in the generated answer $\hat{a}_{i}$ with the compressed context and $L_{\hat{a}}$ is the length of $\hat{a}_{i}$.
@@
where $\beta$ is a hyper-parameter that controls the relative importance of precision and recall. When $\beta$ is 1, the ROUGE-L score is simply the harmonic mean of precision and recall. The greater the $\beta$, the heavier the recall’s weight.
2.4.2 Embedding Inversion
Embedding Inversion (IC) predicts part or all text information based on the embeddings. It is generally regarded as an attacker, with targets including attribute information in the input text, the entire input text, or the text embedding model itself.
@@ 2.4.2 Embedding Inversion
For different levels of leak information, the datasets can be word-level or sentence-level. For attribute inference attack, the dataset has the form of (Input, Words), where the "Input" is text and "words" represents important information in the input text, for example, name and ID number. To reconstruct all the input text, the dataset just contains input text. Note that text embedding will not be contained in the dataset directly; it can be generated by the victim embedding model with the text in the dataset.
@@
The main metrics used for evaluating embedding inversion include BLEU, ROUGE, Exact-Match, cosine similarity, and Token F1. Please refer to Equation 12 and 13 for Exact-Match and ROUGE-L, separately.
  • BLEU is used to measure n-gram similarity between reconstructed text and original input text. we can compute it in the following:
@@
  • Cosine Similarity evaluates the similarity between the embeddings of the reconstructed text and the original input text in the embedding space. It is defined as:
    Cosine $\text{Cosine Similarity}=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$ (17)
    where $\mathbf{a}$ and $\mathbf{b}$ are the embeddings of the reconstructed and original texts, respectively.
  • Token F1 is a word-level metric that computes the F1 score between tokens of the predicted text and the input text. It is defined as the harmonic mean of token-level precision and recall:
[…]
    where BP is the brevity penalty, $w_{n}$ is the weight for each n-gram precision, and $p_{n}$ is the modified precision for n-th n-gram.
Table 3: The Overview of LLM-augmented Text Embedding. I, Q, D+, D-, L denote instruction, query, positive document, hard negative document, and pair similarity separately.
@@
SKICSE [133] LLaMA2 BERT 1M, 276K (Q, D+) $\times$ ZS STS
DenoSent [134] ChatGPT BERT, RoBERTa 1M (Q, D+) $\surd$ ZS STS
@@
AoE [138] ChatGPT, LLaMA2, ChatGLM BERT 6K (Q, D+, D-) $\surd$ ZS STS
@@
SumCSE [140] Vicuna RoBERTa 276K (Q, D+, D-) $\surd$ ZS STS
@@
InPars [147] ChatGPT monoT5 10K (Q, D+, D-) $\surd$ ZS / FS IR
InPars-v2 [148] GPT-J monoT5 180K (Q, D+, D-) $\surd$ ZS / FS IR
NV-Retriever [149] E5mistral Mistral 956K (Q, D+, D-) $\times$ ZS / FT IR
I3 [150] ChatGPT COCO-DR(BERT) 140K (I, Q, D+, D-) $\times$ ZS IR
@@
E5mistral [153] ChatGPT, GPT4 Mistral 500K (I, Q, D+, D-) $\times$ ZS Uni.
@@
3 LLM-Augmented Text Embedding
One approach to adopting LLMs for text embeddings is through knowledge distillation. Specifically, LLMs can be used to generate training data for the embedding model in two ways: 1. Directly synthesizing training data (§ 3.1); 2. Providing supervision signals for existing data (§ 3.2). Please refer to Table 3 for the methods presented in this survey.
3.1 Data Synthesis with LLMs
@@
3.1.1 Instructions
Existing instructions in training data are typically constructed based on datasets [67]. This involves manually collecting multi-task datasets and applying templates to generate different instructions for each dataset [67]. However, this approach lacks diversity, limiting the full potential of the model’s instruction-following capabilities. Therefore, a promising solution is to generate diverse instructions by using LLMs. One approach leverages LLMs to generate various instructions by varying the given conditions: I3 [150] generates diverse instructions by manually setting different conditions, including topics, organizational formats of the retrieved text, and definitions of relevance; E5mistral [153] categorizes tasks into different types (e.g., symmetric and asymmetric retrieval tasks) and designs specific prompts for each category to generate instructions. Another approach leverages LLMs to generate instance-level instructions for each document, further enhancing the diversity of the instructions: Gecko [152] prompts LLMs to generate instructions by conditioning on different documents; Promptriever [151] generates instructions that describe the relationship between the given queries and documents, and enhance instruction diversity by setting various conditions such as length and style of the generated instructions.
3.1.2 Positive Samples
@@
A classic task where anchors and positive samples are symmetric is STS. In this area, numerous studies have proposed various methods to generate symmetric positive samples. Based on the analysis in S-BERT [72], text embedding models for STS are often trained using training data from the Natural Language Inference (NLI) task. Following this convention, a line of research explores how to use LLMs to generate NLI-style training data: NGCSE [143] directly utilize the in-context learning capabilities of LLMs by using NLI-formatted training data as demonstrations to guide the generation of entailment sentences as positive samples; GenSE [135] and MultiCSR [137] first fine-tune LLMs with NLI training data, enabling them to generate NLI-formatted positive samples. Apart from supervised fine-tuning, AdaptCL [141] employ a reinforcement learning approach to iteratively update both the data generation model and the text embedding model, achieving improved performance. In addition, many works focus on exploring generating other types of semantically similar texts as positive samples: SynCSE  [136] and AoE [138] generate diverse semantically similar positive samples by setting different conditions (e.g., genre, topic) and leveraging the in-context learning capabilities of LLMs; SKICSE [133] constructs positive samples by extracting additional information about the anchor using the inherent knowledge of LLMs. SumCSE [140] generates positive samples by having LLMs summarize the anchor; CLAIF [142] creates positive samples by masking tokens in the anchor and completing the missing parts using LLMs; E5mistral [153] further subdivides symmetric retrieval tasks into STS tasks and bitext retrieval tasks, designing specific prompts to generate different types of positive samples for each.
@@
In asymmetric cases, the anchor and positive sample are often referred to as the query and document, respectively. In various studies, either the query or the document can be the target for generation by LLMs. Many works [144, 147, 148, 145, 146, 152] use the vast amount of documents available on the internet to generate queries: Some of them [144, 147, 148, 145, 152] directly leverage the in-context learning capability of LLMs, constructing prompts to facilitate query generation. Notably, InPars [147, 148] introduces a “Guided by Bad Questions (GBQ)” template, which uses randomly sampled pseudo-queries to improve the quality of generated queries; Moreover, SPTAR [146] employs soft prompts to fine-tune LLMs, enabling them to generate queries based on documents. In addition to generating queries from documents, there are other generation methods: I3 [150] and E5mistral [153] generate both queries and documents based on their self-generated instructions; FollowIR [154] and promptriever [151] guides document generation using queries and the generated instructions.
3.1.3 Negative Samples
Negative samples are crucial in contrastive learning, and research [155] has shown that hard negatives, as opposed to randomly sampled negatives from training batches, can significantly improve model performance. Consequently, many studies have explored methods for generating hard negative samples using LLMs. When generating hard negatives, methods similar to those for generating positive samples (as discussed in § 3.1.2) are often employed [135, 137, 139, 141, 143, 136, 138, 140, 151]. The generated hard negatives can generally be divided into two categories. 1. NLI-based negatives [135, 137, 139, 141, 143]: Texts with a logical relationship of contradiction are generated as hard negatives; 2. Contextually irrelevant negatives [136, 138, 140, 151]: Texts are generated using specific definitions of irrelevance as hard negatives. Notably, Promptriever [151] introduces a unique type of hard negative to enhance the model’s instruction-following capability. These hard negatives are positive samples relative to the query itself but become negative samples when the query is combined with the instruction.
3.2 Data Annotation with LLMs